OpenStudy (anonymous):

4x+3y+2z=43 2x+4y+3z=45 3x+2y+4z=47

OpenStudy (anonymous):

i will be bak in 5 min

OpenStudy (anonymous):

k..im back

OpenStudy (anonymous):

now i need help

OpenStudy (anonymous):

wat long replys

OpenStudy (mimi_x3):

x = 13/3 , y=13/3 , z = 19/3 xDD

OpenStudy (mimi_x3):

Probably Saifi is stuck he doesn't know how to do it :P

OpenStudy (anonymous):

howd u get the answers wen u dunt even know how to do them @ H.mini

OpenStudy (anonymous):

lol

OpenStudy (mimi_x3):

LOL, I used my brain and i came up with the answers :P

OpenStudy (mimi_x3):

Hey can you do it like the normal one with only 2 equations >

OpenStudy (anonymous):

well at the back of the book it says (7, 4, 9) answers

OpenStudy (anonymous):

well..i dunno whch is wrong...

OpenStudy (anonymous):

4x+3y+2z=43 [1] 2x+4y+3z=45 [2] 3x+2y+4z=47 [3] To eliminate the z-variables from equations [2] and [3]... Make a linear combination of 3(4x+3y+2z)=3*43 [1]' -2(2x+4y+3z)=-2*45 [2]' ================ [2]'' Now make a linear combo of these -2(4x+3y+2z)=-2*43 [1]'' 3x+2y+4z=47 [3] ============== [3]' Then solve the 2x2 system involving [2]'' and [3]' for x and y then back substitute into one of the original euqations for z. I checked the solutions on wolfram as did mimi, so i won't post them.

OpenStudy (anonymous):

so is wolfram correct?

OpenStudy (mimi_x3):

Well, use the answer you have and substitute it into the equation to see if its correct

OpenStudy (anonymous):

I just checked the wolfram link, and those are the same answers i got; it is not likely we might have made the same typo. wolfram is probably correct; is it possible that you have a typo?

OpenStudy (anonymous):

i didnt get these answers..i got them from the answers at the back of my book and they read (7, 4, 9)

OpenStudy (mimi_x3):

Substitute it back in and see, there might be 2 answers to it

OpenStudy (anonymous):

take mimi's advice it should be easy to check those integer answers in the equations to see if they satisfy

OpenStudy (anonymous):

oh wait..i wrote the question wrong..the first equation is 4x+3y+2z=34

OpenStudy (mimi_x3):

LOL, wonder why

OpenStudy (anonymous):

the ol' typo on the question routine, hmmm...

OpenStudy (anonymous):

so srry...

OpenStudy (mimi_x3):

Let me try it, from what mandolino did xD

OpenStudy (anonymous):

yep..wolfram is correct too wen i tried it..it gave (1 4 9)

OpenStudy (mimi_x3):

Woah~ this is hard, im getting confused xD Saifi got scared that he ran away

OpenStudy (anonymous):

lol..maybe he did..anyways..i got like 30 problems like this...

OpenStudy (mimi_x3):

Sorry I can't help.

OpenStudy (anonymous):

dang it... I HATE MATH!!!!

OpenStudy (mimi_x3):

Me too *high five*

OpenStudy (anonymous):

hero...i made a typo in the first equation its is: 4x+3y+2z=34

hero (hero):

Hmmm. I'm talking to Lana right now, lol

OpenStudy (anonymous):

i dunno who lana is...but i need a good stragety to get through these problems

OpenStudy (anonymous):

wahhh..i need help!!!

OpenStudy (anonymous):

Make a linear combination of 12x+9y+4z=102 [1]' x3 -4x-8y-6z=-90 [2]' x -2 ================ 8x+y =12 [2]'' Now make a linear combo of these -8x-6y-4z=-86 [1]'' x -2 3x+2y+4z=47 [3] ============== -5x-4y =-39 [3]' Now solve the 2x2 system 8x+y =12 [2]'' -5x-4y =-39 [3]' Mult [2]'' by 4 and make a linear combo 32x+4y=48 [2]''' x4 -5x-4y=-39 [3]' ========== 27x =9 x=9/27 = 1/3 I've got a mistake somewhere, but I can't find it. Maybe if I post it someone will find it. It's an arithmetic error. This is a very tedious problem.

hero (hero):

There is no good strategy. You have to do these the hard way

OpenStudy (anonymous):

can u esplain this way

hero (hero):

It's a lot of steps

OpenStudy (anonymous):

i really need help!!!!

hero (hero):

Can you say that with a smile on your face?

OpenStudy (anonymous):

I found the error; I went back to the bloody 43 instead of the 34 in the second linear combo to get [3]' I will redo from there now next post.

OpenStudy (anonymous):

wat post? do i make another?

hero (hero):

bloody math?

OpenStudy (anonymous):

Now make a linear combo of these -8x-6y-4z=-68 [1]'' x -2 3x+2y+4z=47 [3] ============== -5x-4y =-21 [3]' Now solve the 2x2 system 8x+y =12 [2]'' -5x-4y =-21 [3]' Mult [2]'' by 4 and make a linear combo 32x+4y=48 [2]''' x4 -5x-4y=-21 [3]' ========== 27x =27 x=1 There must be another &^%#$% error! since x=13/3

OpenStudy (anonymous):

NOPE..no errors..im jus soooo confused!!!

OpenStudy (anonymous):

Holy &^$$ I am correct; I was using the old answer from wolfram!!!!!! x=1, y=4, z=9

OpenStudy (anonymous):

i still dunt understand ur method...could u kinda esplain it to me? like all the way till every answer?

OpenStudy (anonymous):

Ok. a 3x3 is pretty daunting, so I used elimination to eliminate the same variable (z) from two of the equations [2] and [3]. That's the part where I mutliplied equation [1] and [2] so that the coefficients for z were additive inverses; I did the same thing with [1] and [3]. That's what I called the linear combination. (i will continue on my next post)

OpenStudy (anonymous):

help?

OpenStudy (anonymous):

wat other post?

OpenStudy (anonymous):

the one that i'm writing while you are reading what i just wrote

OpenStudy (anonymous):

oh..

OpenStudy (anonymous):

That gives us the 2x2 system involving equations [2]'' and [3]'. again i decided to use elimination method (but you could just as easily used substitution). in this case i decided to eliminate the y-variable by multiplying [2]'' by 4-- that made the coefficient for y in those two equations opposites so that when I added them together, the y was eliminated and I could solve for x.

OpenStudy (anonymous):

my head hurts...

OpenStudy (anonymous):

At this point, you can sub x=1 into any of the equations in the 2x2 to get y; i choose [2]''... 8(1)+y =12 [2]'' y=12-8=4 Now we have x=1 and y=4

OpenStudy (anonymous):

To get z we have to sub those values for x and y into one of the given equations because it has a z variable in it. i choose equation [3]: 3(1)+2(4)+4z=47 [3] 3+8+4z=47 4z=47-11 4z=36 z=9 Finally we have x=1, y=4, z=9

OpenStudy (anonymous):

i got the last part..trying to get the first 3 posts into my head..

OpenStudy (anonymous):

This was an especially tedious 3x3 problem; I saw that from the beginning which is why I only wanted to set it up, and not solve it. The typo did not help. But these are often not this bad. it is better to learn on a less tedious problem.

OpenStudy (anonymous):

I suggest that you make notes, or copy and paste; then try to redo the problem from the start when you are fresh.

OpenStudy (anonymous):

ya..maybe i will do the rest tmmrw....

OpenStudy (anonymous):

good idea

OpenStudy (anonymous):

oh... thanks for the help hero

OpenStudy (anonymous):

thanks a bunch though...

hero (hero):

Me? What did I do?

OpenStudy (anonymous):

ur welcome :})

OpenStudy (anonymous):

just kidding around, lol

hero (hero):

I would have helped but I was too busy trying to find Lana

OpenStudy (anonymous):

who is this lana?

hero (hero):

I can't tell you

OpenStudy (anonymous):

yea who is lana when she's at home

OpenStudy (anonymous):

oooohhhh!!!!